# MS-Mix: Unveiling the Power of Mixup for Multimodal Sentiment Analysis

### v1

@@ Abstract.
Multimodal Sentiment Analysis (MSA) aims to identify and interpret human emotions by integrating information from heterogeneous data sources such as text, video, and audio. While deep learning models have advanced in network architecture design, they remain heavily limited by scarce multimodal annotated data. Although Mixup-based augmentation improves generalization in unimodal tasks, its direct application to MSA introduces critical challenges: random mixing often amplifies label ambiguity and semantic inconsistency due to the lack of emotion-aware mixing mechanisms. To overcome these issues, we propose MS-Mix, an adaptive, emotion-sensitive augmentation framework that automatically optimizes sample mixing in multimodal settings. The key components of MS-Mix include: (1) a Sentiment-Aware Sample Selection (SASS) strategy that effectively prevents semantic confusion caused by mixing samples with contradictory emotions. (2) a Sentiment Intensity Guided (SIG) module using multi-head self-attention to compute modality-specific mixing ratios dynamically based on their respective emotional intensities. (3) a Sentiment Alignment Loss (SAL) that aligns the prediction distributions across modalities, and incorporates the Kullback-Leibler-based loss as an additional regularization term to train the emotion intensity predictor and the backbone network jointly. Extensive experiments on three benchmark datasets with six state-of-the-art backbones confirm that MS-Mix consistently outperforms existing methods, establishing a new standard for robust multimodal sentiment augmentation. The source code is available at: https://github.com/HongyuZhu-s/MS-Mix.
Multimodal sentiment analysis, Data augmentation, Regularization
copyright: acmlicensedcopyright: acmlicensedjournal: TOMMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Computing methodologies Regularization
1. Introduction
@@
The perception and understanding of human emotions by Artificial Intelligence (AI) is of great significance for technologies such as human-computer interaction (Hibbeln et al., 2017) and multimedia computing(Li et al., 2019). Multimodal Sentiment Analysis (MSA) has emerged as a critical research frontier in this endeavor (Dumpala et al., 2019; Liu et al., 2025; Mao et al., 2022; Zhang et al., 2023; Zhuang et al., 2024; Xiao et al., 2025). MSA aims to integrate and interpret complementary emotional cues from textual, acoustic, and visual modalities to achieve a more robust and accurate understanding of sentiment (Manning et al., 2014). The significance of MSA is underscored by its wide-ranging applications, including but not limited to opinion mining on social media (Chandrasekaran et al., 2021), empathetic chatbot design (Yang et al., 2019), and automated mental health assessment (Singh et al., 2024). By moving beyond unimodal analysis, MSA offers the potential to capture the complex and often incongruent nature of human affective expression, thereby providing a more holistic view of sentiment.
@@
Benefiting from the rapid advances in Deep Learning (DL), the field of MSA has witnessed substantial progress in the development of neural architectures. Techniques such as cross-modal transformers (Tsai et al., 2019) and advanced fusion mechanisms (Zhuang et al., 2024) have demonstrated remarkable capability in modeling inter-modal interactions and extracting discriminative features. Nevertheless, despite these architectural advancements, the performance of data-driven DL models remains fundamentally constrained by the scale and quality of annotated data (Zhang et al., 2018). This data scarcity often results in overfitting and limited generalization ability (Srivastava et al., 2014). Furthermore, the construction of MSA datasets is costly, as it necessitates the collection of reliable human-annotated multimodal sentiment dataset.
@@
Fig. 1. The differences between MS-Mix and traditional mixup methods represented by Manifold Mixup (Zhang et al., 2018). Left: a. The traditional mixup method employs random selection of samples and an offline mixing ratios optimization strategy. b. Backbone. c. MS-Mix. Right: MS-Mix can generate augmented samples that better align with the original distribution on the MOSEI dataset. Where the KL represents the Kullback-Leibler divergence. :
@@
To mitigate the data scarcity issue, the mixup strategy, a data augmentation technique, was introduced to generate virtual training samples via convex interpolations between original data points and their corresponding labels (Zhang et al., 2018). However, existing offline methods, including general domain techniques such as CutMix (Yun et al., 2019) and SaliencyMix (Uddin et al., 2021), as well as multimodal specific approaches like $\mathcal{P}ow$Mix (Georgiou et al., 2024) and MultiMix (Venkataramanan et al., 2023), often depend on randomly paired samples and offline optimization strategies, thereby neglecting the underlying emotional semantics. This may result in the blending of semantically inconsistent samples, e.g., combining a positive and a negative sample can introduce semantic confusion and label mismatch. Moreover, these methods typically employ fixed or uniformly distributed mixing ratios, which fail to adapt to the varied emotional intensities across different modalities. As a consequence, the quality and efficacy of the augmented data are substantially compromised.
@@
To address these limitations, we propose MS-Mix, a novel and adaptive data augmentation framework specifically designed for multimodal sentiment analysis. As shown in Fig. 1, we have visualized the distribution of the generated labels. It can be observed that the label distribution generated by MS-Mix is closer to the real situation. Our method introduces three key innovations to ensure semantically consistent and high-quality sample generation: (1) We design a Sentiment-Aware Sample Selection (SASS) strategy that leverages semantic similarity in the latent space to filter out incompatible pairs, thereby preventing the mixture of samples with contradictory emotions. (2) We introduce a Sentiment Intensity Guided (SIG) mixing module to dynamically determine the mixing ratio across modalities, which leverages a multi-head self-attention mechanism to predict modality-specific mixing ratios conditioned on sentiment salience, thereby enabling more fine-grained and context-aware fusion. (3) We introduce a Sentiment Alignment Loss (SAL) to align the predicted sentiment distribution with the ground-truth labels. The SAL acts as a regularization term, enhancing the backbone network’s discriminative power and improving the consistency of representations from augmented samples. Experimental results conducted on the real-world MSA datasets demonstrate that MS-Mix significantly outperforms existing methods across multiple baselines and backbone architectures.
The main contributions of this paper are summarized as follows:
@@
  • We design an SASS strategy that leverages semantic similarity in the latent space to filter out incompatible sample pairs, effectively preventing mixtures of samples with contradictory emotions.
@@
  • We propose SAL, a Kullback-Leibler (KL) divergence-based regularization that aligns predicted sentiment distributions with ground-truth labels, thereby enhancing both model discrimination and consistency.
  • We conduct extensive experiments on three popular MSA datasets and six backbone architectures, demonstrating that MS-Mix significantly outperforms existing methods.
@@
The remainder of this paper is organized as follows. Section 2 reviews related methods. The proposed MS-Mix method is detailed in Section 3. The experimental results are presented in Section 4, and the concluding remarks are provided in Section 5.
2. Related Work
2.1. Multimodal Sentiment Analysis
@@
Sentiment analysis systems are divided into single-modal systems and multi-modal systems. Compared with single-modal systems, which analyze the sentiment state through a single data source, MSA systems can utilize complementary information from various modalities such as text, video, and audio, thereby providing an effective method for comprehensive sentiment analysis (Geetha et al., 2024). For instance, Cimtay et al. (Cimtay et al., 2020) achieved emotion recognition by extracting data features from Galvanic Skin Response (GSR), Electroencephalogram (EEG), and facial expressions. Prior research in MSA has predominantly advanced through improving two main paradigms: feature fusion strategy-centric (Zadeh et al., 2017; Liu and Shen, 2018; Zadeh et al., 2018a; Tsai et al., 2019; Lv et al., 2021; Zhuang et al., 2024) and feature encoder method-centric (Liang et al., 2021; Han et al., 2021; Yang et al., 2022; Zhang et al., 2023) approaches.
@@
Fusion strategy-centric approaches in MSA primarily focused on designing effective fusion strategies to integrate features from different modalities. Zadeh et al. (Zadeh et al., 2017) introduced the Tensor Fusion Network (TFN), which explicitly models inter-modal interactions through the 3-fold cartesian product. Building on this work, Z Liu et al. (Liu and Shen, 2018) proposed Low-rank Multimodal Fusion (LMF), an efficient approach that enhances computational performance via parallel tensor and weight decomposition. By leveraging modality-specific low-rank factors, LMF achieves effective multimodal fusion with significantly reduced computational cost. MuIT (Tsai et al., 2019) incorporates cross-modal attention, exploring the long-range dependency relationships among cross-modal elements in multimodal datasets, and achieving multimodal fusion on unaligned multimodal datasets. Based on the same motivation and algorithm, Huang et al. (Huang et al., 2020) employed cross-modal attention to learn the long-range dependencies between the visual and audio modalities, achieving better performance than decision-level and feature-level fusion. A recent advancement in MSA is the attempt to decouple the features into shared and unique information. For instance, Li et al. (Li et al., 2023) proposed a Decoupled Multimodal Distillation (DMD) method, which is capable of distilling cross-modal knowledge in the decoupled feature spaces and alleviating the problem of inherent multimodal heterogeneity. The Global Local Modal (GLoMo) (Zhuang et al., 2024) Fusion framework integrates multiple local representations within each modality and effectively combines local and global information.
@@
For feature encoder-centric approaches, the primary objective is to enhance the feature representation of each modality. Yang et al. (Yang et al., 2022) conceptualize multimodal representation learning as a form of domain adaptation, employing adversarial learning to model both modality-invariant and modality-specific subspaces within multimodal fusion. Han et al. (Han et al., 2021) introduced the MMIM framework, which enhances multimodal fusion performance through hierarchical mutual information maximization. To address the challenge of multimodal heterogeneity in MSA, Li et al. (Lin and Hu, 2023) conceptualized the learning process of each modality as a set of sub-tasks and introduced a novel Multi-Task Momentum Distillation (MTMD) method. This approach effectively reduces the discrepancies between modalities and enhances representation consistency. Furthermore, to suppress sentiment-irrelevant and conflicting information across modalities, Zhang et al. (Zhang et al., 2023) proposed the Adaptive Language-guided Multimodal Transformer (ALMT), which learns a unified hypermodality representation guided by language at multiple scales. In recent work, Li et al. (Li et al., 2025) proposed the MM-PEAR-CoT framework, which enhances the reliability of multimodal sentiment analysis by using cross-modal filtering and fusion to suppress irrational reasoning steps generated by large language models.
@@
Despite these advancements, MSA systems remain constrained by the scarcity and high annotation cost of multimodal data, often leading to overfitting and limited generalization (Venkataramanan et al., 2023; Georgiou et al., 2024). This highlights the need for effective data augmentation techniques specifically designed for multimodal emotional data.
2.2. Mixup-based Augmentation
@@
To mitigate data scarcity and enhance generalization, data augmentation techniques—particularly Mixup and its variants—have been extensively adopted in DL (Kim et al., 2020a; Liu et al., 2022; Qin et al., 2024a, b; Jin et al., 2024). The original Mixup method (Zhang et al., 2018) operates in the input space by generating virtual samples through linear interpolation between two randomly selected data points and their corresponding labels. This simple yet effective strategy encourages models to behave linearly across classes and improves robustness. Subsequent extensions such as Manifold Mixup (Verma et al., 2019) generalize the interpolation operation to hidden representations, further enhancing the smoothness of decision boundaries. Similarly, in the latent space, the Mixup-Transformer (Sun et al., 2020) encodes a sentence using a Transformer and then linearly interpolates the representations to generate mixed samples for classification. In recent state-of-the-art approaches (Liu et al., 2022; Qin et al., 2024a, b), mixing strategies are no longer manually designed based on prior knowledge or saliency information but instead adaptively generate mixed samples using learnable mixing ratios and feature representations in an end-to-end manner.
@@
In the multimodal domain, several works have extended Mixup to leverage cross-modal interactions. For instance, MultiMix (Venkataramanan et al., 2023) performs independent Mixup operations within each modality and integrates the augmented representations through late fusion. VLMixer (Wang et al., 2022) integrates cross-modal CutMix (Yun et al., 2019) with contrastive learning to convert uni-modal text inputs into multi-modal representations of text and image, thereby improving instance-level alignment across modalities. To enhance robustness against missing modalities, Lin et al. (Lin and Hu, 2024) proposed Multi-Modal Mixup ($M^{3}$ixup), which extends mixup to both representation and contrastive loss levels for improved cross-modal alignment and dynamics capture. More recently, as an improvement of MultiMix, $\mathcal{P}$owMix (Georgiou et al., 2024) introduced a weight-aware mixing strategy that dynamically modulates mixing coefficients based on the estimated importance of each modality, enabling more flexible and effective multimodal augmentation.
@@
However, most existing methods rely on random sample pairing and offline mixing schemes, which overlook the semantic structure and emotional coherence of multimodal data. This often results in semantically inconsistent mixtures, such as mixing samples with opposing emotions, which introduces label noise and hinders model performance. These limitations underscore the need for more semantically aware and adaptively controlled mixing strategies tailored to multimodal sentiment analysis. Therefore, we propose MS-Mix, a novel mixup method that effectively avoids the mixing of contradictory emotions and adaptively generates discriminative multimodal emotional features.
@@
Fig. 2. An overview of the overall structure of the proposed MS-Mix framework. (a) The SASS strategy computes the emotional semantic distance between samples. (b) The SIG module performs adaptive mixing of sample pairs with similar emotional semantics. (c) The SAL ($\mathcal{L}_{SAL}$) serves as an auxiliary regularization term that promotes alignment between predicted emotional intensity values and ground-truth labels via a KL-based loss. These components work collaboratively to enhance multimodal representation learning. :
3. Methodology

### v2

# MS-Mix: An Adaptive Emotion-Sensitive Mixup Framework for Multimodal Sentiment Analysis

@@ Abstract
Multimodal Sentiment Analysis (MSA) integrates complementary features from text, video, and audio to achieve robust emotion understanding. Yet, models for this task suffer from severe data scarcity and high annotation costs. This severely restricts their practical engineering applications. Although Mixup-based augmentation improves generalization in unimodal settings, its naive application to MSA often produces semantically inconsistent mixtures and amplified label noise due to random sample mixing and fixed mixing ratios that ignore emotional semantics. To overcome these issues, we propose MS-Mix, an adaptive, emotion-sensitive augmentation framework that automatically optimizes sample mixing in multimodal settings. The key components of MS-Mix include: (1) a Sentiment-Aware Sample Selection strategy that filters incompatible pairs by measuring semantic similarity in the latent space, preventing mixtures of contradictory emotions. (2) a Sentiment Intensity Guided module using multi-head self-attention, that dynamically computes modality-specific mixing ratios conditioned on emotional salience. (3) a Sentiment Alignment Loss based on Kullback-Leibler divergence that regularizes the model by aligning predicted sentiment distributions across modalities with ground-truth labels, improving discrimination and consistency. Extensive experiments on three benchmark datasets with six state-of-the-art backbones confirm that MS-Mix consistently outperforms existing methods, thereby improving the robustness and practical applicability of MSA in engineering contexts. The source code is available at an anonymous link: https://anonymous.4open.science/r/MS-Mix-review-0C72.
keywords:
Affective computing , Multimodal sentiment analysis ,
Data augmentation , Regularization
journal: Nuclear Physics B
[…]
[1]organization=Chongqing Institute of Green Intelligent Technology, Chinese Academy of Sciences; Chongqing School, University of Chinese Academy of Sciences, addressline=400714, city=Chongqing, country=China
[…]
[2]organization=School of Engineering, Westlake University, city=Zhejiang, country=China
[…]
[3]organization=Telecom SudParis, Institute Polytechnique de Paris, city=Palaiseau, country=France
1 Introduction
@@
The perception and understanding of human emotions by Artificial Intelligence (AI) are of great significance for technologies such as human-computer interaction [1], and multimedia computing [2]. Multimodal Sentiment Analysis (MSA) has emerged as a critical research frontier in this endeavor [3, 4, 5, 6, 7]. MSA aims to integrate and interpret complementary emotional features from textual, acoustic, and visual modalities to achieve a more robust and accurate understanding of sentiment [8]. The significance of MSA is underscored by its wide-ranging applications, including but not limited to opinion mining on social media [9], empathetic chatbot design [10], and mental health assessment [11]. By moving beyond unimodal analysis, MSA can capture the complex, often incongruent nature of human affective expression, thereby providing a more holistic view of sentiment.
@@
Benefiting from rapid advances in Deep Learning (DL), the field of MSA has witnessed substantial progress in developing neural architectures. Techniques such as cross-modal transformers [12] and advanced fusion mechanisms [6] have demonstrated remarkable capability in modeling inter-modal interactions and extracting discriminative features. Nevertheless, despite these architectural advancements, the performance of data-driven DL models remains fundamentally constrained by the scale and quality of annotated data [13]. This data scarcity often results in overfitting and limited generalization ability [14]. Furthermore, constructing MSA datasets is costly, as it requires the collection of reliable, human-annotated multimodal sentiment datasets. Above all, these issues severely restrict the practical engineering applications of MSA.
@@
Figure 1: The differences between MS-Mix and traditional mixup methods represented by Manifold Mixup [13]. Left: a. The traditional methods employ random sampling and an offline optimization of mixing ratios. b. Backbone. c. MS-Mix. Right: MS-Mix can generate augmented samples that better align with the original label distribution.
@@
To mitigate data scarcity, the mixup strategy, a data augmentation technique, was introduced to generate virtual training samples by convexly interpolating between original data points and their corresponding labels [13]. However, existing offline methods, including general domain techniques such as CutMix [15] and SaliencyMix [16], as well as multimodal specific approaches like $\mathcal{P}ow$Mix [17] and MultiMix [18], often depend on randomly paired samples and offline optimization strategies, thereby neglecting the underlying emotional semantics. This may result in the blending of semantically inconsistent samples, e.g., combining a positive and a negative sample can introduce semantic confusion and label mismatch. Moreover, these methods typically use fixed or uniformly distributed mixing ratios, which do not adapt to the varying emotional intensities across modalities. As a consequence, the quality and efficacy of the augmented data are substantially compromised.
@@
To address these limitations, we propose MS-Mix, a novel, adaptive data augmentation framework designed for multimodal sentiment analysis. As shown in Fig. 1, we have visualized the distribution of the generated labels, and MS-Mix has a lower Kullback- Leibler (KL) divergence. It can be observed that the label distribution generated by MS-Mix is closer to the real situation. Our method introduces three key innovations to ensure semantically consistent and high-quality sample generation: (1) We design a Sentiment-Aware Sample Selection (SASS) strategy that leverages semantic similarity in the latent space to filter out incompatible pairs, thereby preventing the mixture of samples with contradictory emotions. (2) We introduce a Sentiment Intensity Guided (SIG) mixing module to dynamically determine the mixing ratio across modalities, which leverages a multi-head self-attention mechanism to predict modality-specific mixing ratios conditioned on sentiment salience, thereby enabling more fine-grained and context-aware fusion. (3) We introduce a Sentiment Alignment Loss (SAL) to align the predicted sentiment distribution with the ground-truth labels. The SAL acts as a regularization term, enhancing the backbone network’s discriminative power and improving the consistency of representations from augmented samples. Experimental results on real-world MSA datasets show that MS-Mix significantly outperforms existing methods across multiple baselines and backbone architectures. The main contributions of this paper are summarized as follows:
  • 1.
@@
    We design a SASS strategy that leverages semantic similarity in the latent space to filter out incompatible sample pairs, effectively preventing mixtures of samples with contradictory emotions.
  • 2.
@@
    We propose SAL, a Kullback-Leibler divergence-based regularization that aligns predicted sentiment distributions with ground-truth labels, thereby enhancing both model discrimination and consistency.
@@
The remainder of this paper is organized as follows: Section 2 reviews related works. The proposed MS-Mix method is detailed in Section 3. The experimental results and visualizations are presented in Section 4, and the conclusions are provided in Section 5.
2 Related Work
2.1 Multimodal Sentiment Analysis
@@
Sentiment analysis systems are divided into single-modal systems and multimodal systems. Compared with single-modal systems, which analyze sentiment using a single data source, MSA systems can leverage complementary information across multiple modalities, such as text, video, and audio, thereby providing a more comprehensive approach to sentiment analysis [19]. For instance, Cimtay et al. [20] achieved emotion recognition by extracting data features from Galvanic Skin Response (GSR), Electroencephalogram (EEG), and facial expressions. Prior research in MSA has predominantly focused on improving two main paradigms: feature-fusion strategy-centric [21, 22, 23, 12, 24, 6] and feature-encoder method-centric [25, 26, 27, 5] approaches.
@@
Fusion strategy-centric approaches in MSA primarily focused on designing effective fusion strategies to integrate features from different modalities. Zadeh et al. [21] introduced the Tensor Fusion Network (TFN), which explicitly models intermodal interactions via the 3-fold Cartesian product. Building on this work, Z. Liu et al. [22] proposed Low-rank Multimodal Fusion (LMF), an efficient approach that improves computational performance through parallel tensor and weight decomposition. By leveraging modality-specific low-rank factors, LMF achieves effective multimodal fusion with significantly reduced computational cost. MuIT [12] incorporates cross-modal attention, explores long-range dependencies among cross-modal elements in multimodal datasets, and achieves multimodal fusion across unaligned datasets. A recent advancement in MSA is the attempt to decouple the features into shared and unique information. For instance, Li et al. [28] proposed a Decoupled Multimodal Distillation (DMD) method that distills cross-modal knowledge in decoupled feature spaces and alleviates the inherent multimodal heterogeneity. The Global Local Modal (GLoMo) [6] Fusion framework integrates multiple local representations within each modality and effectively combines local and global information.
@@
In feature encoder-centric approaches, the primary objective is to improve the feature representation for each modality. Yang et al. [26] conceptualize multimodal representation learning as a form of domain adaptation, employing adversarial learning to model both modality-invariant and modality-specific subspaces within multimodal fusion. To address the challenge of multimodal heterogeneity in MSA, Li et al. [27] conceptualized the learning process of each modality as a set of sub-tasks and introduced a novel Multi-Task Momentum Distillation (MTMD) method. This approach effectively reduces the discrepancies between modalities and enhances representation consistency. Furthermore, to suppress sentiment-irrelevant and conflicting information across modalities, in recent work, Zhang et al. [5] proposed the Adaptive Language-guided Multimodal Transformer (ALMT), which learns a unified hypermodality representation guided by language at multiple scales.
@@
Despite these advancements, MSA systems remain constrained by the scarcity and high annotation costs of multimodal data, which often lead to overfitting and limited generalization [18, 17]. This highlights the need for effective data augmentation techniques specifically designed for multimodal emotional data.
2.2 Mixup-based Augmentation
@@
To mitigate data scarcity and improve generalisation, data augmentation techniques, particularly Mixup and its variants, have been widely adopted in DL [29, 30, 31, 32, 33]. The original Mixup method [13] operates in the input space by generating virtual samples through linear interpolation between two randomly selected data points and their corresponding labels. This simple yet effective strategy encourages models to behave linearly across classes and improves robustness. Subsequent extensions, such as Manifold Mixup [34], generalize the interpolation operation to hidden representations, further enhancing the smoothness of decision boundaries. Similarly, in the latent space, the Mixup-Transformer [35] encodes a sentence using a Transformer and then linearly interpolates the representations to generate mixed samples for classification. In the latest work, AdvMixUp [33] introduces a sample-dependent, feature-level interpolation mask generated by a compact Mask Generator, optimized via adversarial training (min-max framework) to produce challenging ”hard” mixed samples near decision boundaries. Recent state-of-the-art approaches [30, 31, 33] that use mixing strategies no longer rely on manually designed rules based on prior knowledge or saliency information, but instead adaptively generate mixed samples using learnable mixing ratios and feature representations in an end-to-end manner.
@@
In the multimodal domain, several works have extended Mixup to leverage cross-modal interactions. For instance, MultiMix [18] performs independent Mixup operations within each modality and integrates the augmented representations through late fusion. VLMixer [36] integrates cross-modal CutMix [15] with contrastive learning to convert uni-modal text inputs into multi-modal representations of text and image, thereby improving instance-level alignment across modalities. To enhance robustness against missing modalities, Lin et al. [37] proposed Multi-Modal Mixup ($M^{3}$ixup), which extends mixup to both representation and contrastive loss levels to improve cross-modal alignment and capture of dynamics. More recently, as an improvement of MultiMix, $\mathcal{P}$owMix [17] introduced a weight-aware mixing strategy that dynamically modulates mixing coefficients based on the estimated importance of each modality, enabling more flexible and effective multimodal augmentation.
@@
However, most existing methods rely on random sampling and offline mixing, which overlook the semantic structure and emotional coherence of multimodal data. This often results in semantically inconsistent mixtures, such as mixing samples with opposing emotions, which introduces label noise and hinders model performance. These limitations underscore the need for more semantically aware and adaptively controlled mixing strategies tailored to multimodal sentiment analysis. Therefore, we propose MS-Mix, a novel mixup method that effectively avoids the mixing of contradictory emotions and adaptively generates discriminative multimodal emotional features.
@@
Figure 2: An overview of the proposed MS-Mix framework. (a) The SASS strategy computes the emotional semantic distance between samples. (b) The SIG module performs adaptive mixing of sample pairs with similar emotional semantics. (c) The SAL ($\mathcal{L}_{SAL}$) serves as an auxiliary regularization term that promotes alignment between predicted emotional intensity values and ground-truth labels via a KL-based loss. These components work collaboratively to enhance multimodal representation learning.
3 Methodology